# Five-Term Relations for wreath Macdonald polynomials and tableau formulas for Pieri coefficients

### v1

@@ 7.5. Another evaluation formula
For any positive integers $k,l$, and $\mu\subset_{k+l}\lambda$,
${c^{\prime}}_{\mu,\lambda}^{(p)}=\frac{(-1)^{l}}{e_{l}\left[B_{\mu/\lambda}^{(s)}\right]}\sum_{\lambda\subset_{k}\nu\subset_{l}\mu}{c^{\prime}}_{\lambda,\nu}^{(p)}{c^{\prime}}_{\nu,\mu}^{(p)}T_{\mu/\nu}^{(s)}.$
@@ 8.1. Wreath Theta operators
$\left[\Theta^{(s)}_{\tilde{e}^{(p)}_{n}},\mathbb{D}_{\epsilon_{0}-\epsilon_{i}-\epsilon_{s}}\right]=\sum_{k=1}^{n}(-1)^{k}\mathbb{D}_{\epsilon_{0}-\epsilon_{i}-\epsilon_{s}-k\epsilon_{p}}\Theta^{(s)}_{\tilde{e}^{(p)}_{n-k}}.$
@@
$\Theta^{(p,s)}(w;v)\mathbb{D}_{\epsilon_{0}-\epsilon_{i}-\epsilon_{s}}\left(\Theta^{(p,s)}(w;v)\right)^{-1}$ $=\left(\Delta_{v}^{(s)}\mathcal{P}^{(p)}_{-{w}/{M^{T}}}\Delta_{-v}^{(s)}\right)\nabla^{(s)}\tilde{e}_{1}^{(i)}{\nabla^{(s)}}^{-1}\left(\Delta_{v}^{(s)}\mathcal{P}^{(p)}_{w/M^{T}}\Delta_{-v}^{(s)}\right)$
@@
$=\mathcal{P}_{-w/M^{T}}^{(p)}\mathbb{D}_{\epsilon_{0}-\epsilon_{i}-\epsilon_{s}}\mathcal{P}_{w/M^{T}}^{(p)}$
$=\mathcal{P}_{-w/M^{T}}^{(p)}(z_{0}/z_{i}z_{s})\mathbb{D}\mathcal{P}_{w/M^{T}}^{(p)}\Big{|}_{z^{0}}$
$=(1-w/z_{p})^{-1}(z_{0}/z_{i}z_{s})\mathbb{D}\Big{|}_{z^{0}}$
$=\sum_{k\geq 0}w^{k}\mathbb{D}_{\epsilon_{0}-\epsilon_{i}-\epsilon_{s}-k\epsilon_{p}}.$
@@
$\Theta^{(p,s)}(w;v)\mathbb{D}_{\epsilon_{0}-\epsilon_{i}-\epsilon_{s}}=\sum_{k\geq 0}w^{k}\mathbb{D}_{\epsilon_{0}-\epsilon_{i}-\epsilon_{s}-k\epsilon_{p}}\Theta^{(p,s)}(w;v)$
@@
$\Theta^{(s)}_{\tilde{e}^{(p)}_{n}}\mathbb{D}_{\epsilon_{0}-\epsilon_{i}-\epsilon_{s}}=\sum_{k=0}^{n}(-1)^{k}\mathbb{D}_{\epsilon_{0}-\epsilon_{i}-\epsilon_{s}-k\epsilon_{p}}\Theta^{(s)}_{\tilde{e}^{(p)}_{n-k}}.\qed$
@@ A.1. From Pieri to Delta
$d_{\mu,\lambda}^{(p)}=\frac{T^{(0)}_{\lambda}}{T_{\mu}^{(0)}}\mathfrak{d}_{\mu,\lambda}=\chi_{0}^{-1}\mathfrak{d}_{\mu,\lambda}.$
@@ A.1.1. Applying $\mathbb{D}$
$e_{1}\left[\epsilon_{-p}\frac{D^{\dagger}}{M}\right]=\mathbb{D}_{\epsilon_{0}-\epsilon_{p}}.$
@@ A.1.4. Calculation of coefficients
$\mathfrak{d}_{\mu,\lambda}^{(p)}$ $=\frac{1}{1-qt}\left(\frac{\chi_{0}^{2}}{\chi_{p}}\right)\prod_{i\in\mathbb{Z}/r\mathbb{Z}}\frac{\displaystyle\prod_{\square\in R_{i}(\mu)}(\chi_{i}-qt\chi_{\square})}{\displaystyle\prod_{\square\in A_{i}(\mu)}(\chi_{i}-\chi_{\square})}$
@@
(cf. the proof of [GT96, Theorem 2.2]). We obtain (A.1) by dividing by $\chi_{0}$ to obtain $d_{\mu,\lambda}^{(p)}$.

### v3

@@ 7.5. Another evaluation formula
For any positive integers $k,l$, and $\lambda\subset_{k+l}\mu$,
${c^{\prime}}_{\lambda,\mu}^{(p)}=\frac{(-1)^{l}}{e_{l}\left[B_{\mu/\lambda}^{(s)}\right]}\sum_{\lambda\subset_{k}\nu\subset_{l}\mu}{c^{\prime}}_{\lambda,\nu}^{(p)}{c^{\prime}}_{\nu,\mu}^{(p)}T_{\mu/\nu}^{(s)}.$
@@ 8.1. Wreath Theta operators
$\left[\Theta^{(s)}_{\tilde{e}^{(p)}_{n}},\mathbb{D}_{\epsilon_{s}-\epsilon_{0}-\epsilon_{i}}\right]=\sum_{k=1}^{n}(-1)^{k}\mathbb{D}_{\epsilon_{s}-\epsilon_{0}-\epsilon_{i}-k\epsilon_{p}}\Theta^{(s)}_{\tilde{e}^{(p)}_{n-k}}.$
@@
$\Theta^{(p,s)}(w;v)\mathbb{D}_{\epsilon_{s}-\epsilon_{0}-\epsilon_{i}}\left(\Theta^{(p,s)}(w;v)\right)^{-1}$ $=\left(\Delta_{v}^{(s)}\mathcal{P}^{(p)}_{-{w}/{M^{T}}}\Delta_{-v}^{(s)}\right)\nabla^{(s)}\tilde{e}_{1}^{(i)}{\nabla^{(s)}}^{-1}\left(\Delta_{v}^{(s)}\mathcal{P}^{(p)}_{w/M^{T}}\Delta_{-v}^{(s)}\right)$
@@
$=\mathcal{P}_{-w/M^{T}}^{(p)}\mathbb{D}_{\epsilon_{s}-\epsilon_{0}-\epsilon_{i}}\mathcal{P}_{w/M^{T}}^{(p)}$
$=\mathcal{P}_{-w/M^{T}}^{(p)}(z_{s}/z_{0}z_{i})\mathbb{D}\mathcal{P}_{w/M^{T}}^{(p)}\Big{|}_{z^{0}}$
$=(1-w/z_{p})^{-1}(z_{s}/z_{0}z_{i})\mathbb{D}\Big{|}_{z^{0}}$
$=\sum_{k\geq 0}w^{k}\mathbb{D}_{\epsilon_{s}-\epsilon_{0}-\epsilon_{i}-k\epsilon_{p}}.$
@@
$\Theta^{(p,s)}(w;v)\mathbb{D}_{\epsilon_{s}-\epsilon_{0}-\epsilon_{i}}=\sum_{k\geq 0}w^{k}\mathbb{D}_{\epsilon_{s}-\epsilon_{0}-\epsilon_{i}-k\epsilon_{p}}\Theta^{(p,s)}(w;v)$
@@
$\Theta^{(s)}_{\tilde{e}^{(p)}_{n}}\mathbb{D}_{\epsilon_{s}-\epsilon_{0}-\epsilon_{i}}=\sum_{k=0}^{n}(-1)^{k}\mathbb{D}_{\epsilon_{s}-\epsilon_{0}-\epsilon_{i}-k\epsilon_{p}}\Theta^{(s)}_{\tilde{e}^{(p)}_{n-k}}.\qed$
@@ A.1. From Pieri to Delta
$d_{\mu,\lambda}^{(p)}=\frac{T^{(0)}_{\lambda}}{T_{\mu}^{(0)}}\mathfrak{d}_{\mu,\lambda}=-\chi_{0}^{-1}\mathfrak{d}_{\mu,\lambda}.$
@@ A.1.1. Applying $\mathbb{D}$
$-e_{1}\left[\epsilon_{-p}\frac{D^{\dagger}}{M}\right]=\mathbb{D}_{\epsilon_{0}-\epsilon_{p}}.$
@@ A.1.4. Calculation of coefficients
$-\mathfrak{d}_{\mu,\lambda}^{(p)}$ $=\frac{1}{1-qt}\left(\frac{\chi_{0}^{2}}{\chi_{p}}\right)\prod_{i\in\mathbb{Z}/r\mathbb{Z}}\frac{\displaystyle\prod_{\square\in R_{i}(\mu)}(\chi_{i}-qt\chi_{\square})}{\displaystyle\prod_{\square\in A_{i}(\mu)}(\chi_{i}-\chi_{\square})}$
@@
(cf. the proof of [GT96, Theorem 2.2]). We obtain (A.1) by dividing by $-\chi_{0}$ to obtain $d_{\mu,\lambda}^{(p)}$.